OpenStudy (anonymous):

what is the slope-intercept form of this equation? 3x+9y=18

OpenStudy (anonymous):

The equation for the slope-intercept form of a line is y = mx + b where m = slope and b = y intercept. So basically, we just need to manipulate this question a little so we get y = something See what you can do, let me know if you get stuck.

OpenStudy (anonymous):

well that didn't answer my question but thanks....I guess

OpenStudy (anonymous):

Your welcome....

OpenStudy (anonymous):

3x+9y=18 -3x -3x 9y=-3x+18 /9 /9 y=-3/9x+2 simplify y=-1/3x+2